# When Large Vision-Language Model Meets Large Remote Sensing Imagery: Coarse-to-Fine Text-Guided Token Pruning

### v2

@@ 1 Introduction
Benefiting from the rapid advancement of Large Language Models (LLMs) [1, 9, 15, 67, 12], Large Visual-Language Models (LVLMs) have demonstrated strong capabilities in perceiving and understanding visual information in text-based multimodal interactions  [35, 25, 57, 7]. Currently, LVLMs have been extensively studied and applied across various fields like remote sensing (RS) intelligent interpretation [46, 79, 75, 65, 28, 10, 43].
Figure 1: High-resolution strategy comparison for modular LVLMs. (a) and (b) show existing grid-based cropping methods face challenges when processing large RSIs. (c) The proposed dynamic pyramid-based token pruning strategy can dynamically select image tiles of key regions related to the input text, balancing image detail and computational cost.
Advances in satellite imaging technology allow for the acquisition of large remote sensing images (RSIs) that cover extensive ground areas and contain detailed land use information. Therefore, natural language-based intelligent analysis of large RSIs is valuable for applications such as complex scene understanding [33], urban planning [56, 31], infrastructure development [62], and monitoring [85].
@@ 2.2 Vision Token Reduction in LVLMs
Recently, a variety of token pruning methods have been proposed for LVLMs, which can be categorized into three main types: token pruning conducted in the vision encoder  [70, 50, 29, 4, 59, 69], token pruning performed in the LLM component [6, 64, 73, 17, 80, 18], and collaborative approaches that integrate both [26, 77, 5, 27, 32, 66, 74].
@@ 3 Preliminaries
In this section, we present the foundational concepts of text-image attention within the LLM component of high-resolution LVLMs. The proposed RFM leverages this mechanism to guide the text-related key image tile selection and the vision token pruning.
Mainstream modular LVLMs include three components [41]: a pre-trained vision encoder, a feature projector, and a pre-trained decoder-only LLM. Conventional high-resolution strategies [8, 16] process the original image $I_{img}$ into a thumbnail view $I_{\text{thumb}}$ and a set of image tiles $I_{\text{tiles}}$. Then the vision encoder and the projector transforms $I_{\text{thumb}}$ and $I_{\text{tiles}}$ into vision token embeddings $T_{vis}=[T^{lr}_{vis},T^{hr}_{vis}]$, which contain low-resolution tokens $T^{lr}_{vis}$ from the $I_{\text{thumb}}$ and higher-resolution $T^{hr}_{vis}$ from the $I_{\text{tiles}}$. $T_{vis}$ are concatenated with text token embeddings $T_{txt}$ from the tokenized text instruction, forming a multimodal token sequence $T=[T_{vis},T_{txt}]$. The LVLM processes $T$ via transformer-based decoder layers with causal self-attention, where the attention score $A$ can be represented as:
@@
Figure 3: The proposed RFM and attention distillation strategy. The left part indicates our core idea: distill accurate text-related key region localization ability from the LLM part of the LVLM. The right part shows the distillation details. We only select specific layer pairs for distillation to avoid hidden state discontinuities. “sys token” represents the tokens from the system prompt.
@@ 4.1 Method
To achieve efficient high-resolution image understanding, we propose a method that combines text-guided key region localization with a multi-level dynamic image pyramid (DIP), as shown in Fig. 2. Initially, the DIP is constructed based on the input large image. At the low-resolution DIP level, the Region Focus Module (RFM) provides attention distribution for the initial vision tokens. These result guides the retrieval of corresponding image tiles from higher-resolution DIP levels or trigger token pruning at the current level. This iterative process could continue through the pyramid until reaching the original resolution. Below, we introduce our approach in three components: DIP construction in Sec. 4.1.1, RFM and attention distillation in Sec. 4.1.2, and their integration in Sec. 4.1.3.
@@ 4.1.1 The Construction of Dynamic Image Pyramid
Given an original large RSI $I_{img}$, we construct an image-level pyramid with a variable number of layers to generate a series of image tiles at different ground sample distances (GSDs), enabling coarse-to-fine inference.
First, we iteratively downsampling $I_{img}$ by a factor of 2, generating a series of images $\{I^{1}_{init},I^{2}_{init},\dots,I^{P}_{init}\}$ until the shorter side reaches a pre-defined minimum length. Let a basic image tile size be $B\times B$ (e.g., $336\times 336$ for CLIP-L14), for the $p$-th scaled image $I^{p}_{init}$, we calculate the number of image tiles along the height and width as follows:
@@ 4.1.2 Attention Distillation for Key Region Focus
In this section, we introduce the core idea and details of the Region Focus Module (RFM), a lightweight module designed to replicate the text-guided region localization capability inherent in LVLM’s LLM part. RFM can be seamlessly integrated during the supervised fine-tuning (SFT) stage of LVLM, and generates attention scores for vision tokens during inference.
Prior findings [20, 64, 73] demonstrate that the deep layers in the LLM part of LVLM could accurately localize text-related key regions in the image through cross-model attention. Inspired by this observation, our core idea is to distill this text-guided key region focus capability from the LLM, as shown in the left part of Fig. 3. After distillation, the trained RFM can help to select text-related key image tiles and prune vision tokens before the LLM, reducing computational overhead from the image encoding stage.
Specifically, assuming the LLM has $M$ layers $l_{1},l_{2},\dots,l_{M}$, the proposed RFM adopts the same architecture with $R$ layers $r_{1},r_{2},\dots,r_{R}$, where $R$$<$$M$. Each RFM layer $r_{i}$ is initialized from a selected LLM layer $l_{m_{i}}$, where $\{m_{1},m_{2},\dots,m_{R}\}$ forms a sparse subset of LLM layer indices. For layer-wise distillation, we select $K$ pairs ($K$$<$$R$) from RFM and LLM as student-teacher pairs, represented as $\{(r_{k},l_{m_{k}})\mid k=1,2,\dots,K\}$. This strategy avoids rigidly aligning each $r_{i}$ with $l_{m_{i}}$, which would neglect the non-teacher layers in the LLM, resulting in inconsistency between adjacent RFM layers when training.
@@ 4.1.3 Text-Guided Token Pruning with Pyramid
Given the extensive irrelevant background content in large RSIs, the core of our strategy lies in iteratively selecting text-related key image tiles and performing token pruning, thereby achieving a coarse-to-fine inference, as illustrated in the below part of Fig. 2.
Specifically, we integrate DIP and RFM for inference, as detailed in Alg. 1. For simplicity, the tokens from the system prompt are omitted. We initialize the vision input using $I_{\text{thumb}}$ and $I^{p}_{\text{tiles}}$ to generate the initial $T_{vis}=[T^{lr}_{vis},T^{p,hr}_{vis}]$, where $p=1$, then the concatenated tokens $T$ are fed into the RFM. From the last layer of the RFM, we compute the average attention scores from all heads in the MHSA, and extract the attention scores $\smash{A^{K,hr}_{\text{stu}}}$ of $\smash{T^{p,hr}_{\text{vis}}}$. By selecting the top-$\lambda$ proportion from $\smash{A^{K,hr}_{\text{stu}}}$, we identify the indices of key tokens and map their coordinates to image tile-level coordinates, to select key image tiles $I^{p+1}_{\text{key}}$ from $I^{p+1}_{\text{tiles}}$. Based on the number of $I^{p+1}_{\text{key}}$, we determine whether to prune the $\smash{T^{p,hr}_{\text{vis}}}$ directly or replace it with vision tokens $\smash{T^{p+1,hr}_{\text{vis}}}$ from higher-resolution tiles. If $\smash{T^{p+1,hr}_{\text{vis}}}$ is obtained, the tile selection process is repeated recursively until the last layer of DIP is reached. This approach enables LVLM to focus only on processing a few high-resolution image tiles in a coarse-to-fine manner, thereby reducing computational complexity while preserving critical text-related image details.
@@ 4.2 The Construction of LRS-VQA Benchmark
Considering the limitations of existing benchmarks for evaluating LVLMs’ perception of RSIs in terms of question diversity and image size, we propose a new benchmark for more comprehensively evaluation.
@@ 4.2.1 Annotation Pipeline
The annotation pipeline of LRS-VQA is illustrated in Fig. 4. We collect 3 publicly remote sensing datasets: FAIR1M-1.0 [54], GLH-Bridge [31], and STAR [33], and identify all unique targets based on the object detection labels to generate unique references from them (e.g., “the top-most ship” and “the largest dock”). Subsequently, we crop the region with the unique target from the original image, and feed it, along with a prompt containing the unique reference, into GPT-4V [47] to generate question-answer pairs about the target’s color, shape, status, etc. Details of the overall construction process are provided in the Appendix A.1.
@@ 5 Experiments
w/ PruMerge++ 43.27 28.38 32.78 34.86 0.152
w/ VisionZip 42.71 24.55 34.55 33.98 0.183
w/ FastV 39.52 30.10 43.99 37.93 0.192
w/ PDrop 41.00 30.54 47.77 39.85 0.184
@@
Table 3: Comparison with different token reduction methods based on LLaVA-Next-Qwen2. To enable a fair comparison, we adopt the “anyres-p25” setting, where the maximum number of pre-defined grids is 25 (1,680×1,680 pixels). FPS is computed based on the inference time for 800 images from MME-Realworld-RS.
@@
w/ PruMerge++ 83520 5328 43.75
w/ VisionZip 83520 9280 83.56
w/ FastV 83520 21312 109.21
w/ PDrop 83520 21312 101.62
@@ 5.1 Experimental Details
Data. During the pre-training (PT) phase, we utilized the same 558K data used in LLaVA-1.5 [36]. For the SFT phase, we collect 484K samples, including 300K sampled from LLaVA-1.5-665K, 146K sampled from RSVQA-HR [40], and 38K template-based samples from 3 RS datasets [13, 31, 33] using their labels. Note some images in this 38k data are from the same source as LRS-VQA, but there is no overlap or duplication between them. Details can be found in the Appendix A.2.1.
Benchmarks and evaluation metrics: i) MME-RealWorld-RS: the RS part of MME-Realworld [83], containing 1,298 RSIs with expert-annotated questions in three types: color, count, and position. We follow the official evaluation script but modify the prompt by removing “The best answer is:” to address Vicuna-1.5-based models’ tendency to respond with only option A. ii) LRS-VQA: it consists of 3 parts: LRS-FAIR, LRS-Bridge, and LRS-STAR, containing 2,272, 1,062, and 3,999 QA pairs, respectively. For the short open-ended format, we adopt a structured evaluation metric following [19, 63], using WordNet [45], with a semantic similarity threshold of 0.8.
Experimental settings. We employ existing LVLMs’ official weights or api for evaluation. For comparison, we reproduce other methods using the same 484k SFT data. Since the LLaVA-NeXT’s training code is not open-source, we integrate the anyres strategy and Qwen2 into the LLaVA-1.5 framework, which serves as the codebase for our approach. All the experiments are conducted on 4 NVIDIA A100 80GB GPUs. For our method, the minimum length for DIP is 1,008 pixels. For Vicuna-1.5 and Qwen2, the $N_{\max}$ is set to 40 and 80, respectively. The vision token saving ratio $\alpha$ is 0.25, $\lambda_{\text{hr}},\lambda_{\text{mse}},\lambda_{\text{kl}}$ are 2.0, 1.0, 1.0, respectively. The RFM-LLM layer pairs are $[1,5,11,14]$ with distillation applied to the first and last pairs. More experimental details are in the Appendix A.2.2.

### v3

@@ 1 Introduction
Benefiting from the rapid advancement of Large Language Models (LLMs) [2, 11, 17, 75, 14], Large Visual-Language Models (LVLMs) have demonstrated strong capabilities in perceiving and understanding visual information in text-based multimodal interactions  [41, 29, 65, 9]. Currently, LVLMs have been extensively studied and applied across various fields like remote sensing (RS) intelligent interpretation [53, 87, 94, 73, 32, 12].
Figure 1: High-resolution strategy comparison for modular LVLMs. (a) and (b) show that existing grid-based cropping methods face challenges when processing large RSIs. (c) The proposed dynamic pyramid-based token pruning strategy can dynamically select image tiles of key regions related to the input text, balancing image detail and computational cost.
Advances in satellite imaging technology allow for the acquisition of large remote sensing images (RSIs) that cover extensive ground areas and contain detailed land use information [18, 38]. Therefore, language-based intelligent analysis of large RSIs is valuable for applications such as complex scene understanding [37, 50], urban planning [64, 35], infrastructure development [70], and monitoring [93].
@@ 2.2 Vision Token Reduction in LVLMs
Recently, a variety of token pruning methods have been proposed for LVLMs, which can be categorized into three main types: token pruning conducted in the vision encoder  [78, 58, 33, 5, 67, 77], token pruning performed in the LLM component [8, 72, 81, 20, 88, 21], and collaborative approaches that integrate both [30, 85, 6, 31, 36, 74, 82]. Moreover, adaptive token segmentation methods [28, 1, 7] also offer a promising approach for token reduction.
@@ 3 Preliminaries
In this section, we present the foundational concepts of text-image attention within the LLM component of LVLMs. Mainstream modular LVLMs include three components [48]: a pre-trained vision encoder, a projector, and a pre-trained decoder-only LLM. Conventional high-resolution strategies [10, 19] process the original image $I_{img}$ into a thumbnail view $I_{\text{thumb}}$ and a set of image tiles $I_{\text{tiles}}$. Then the vision encoder and the projector transforms $I_{\text{thumb}}$ and $I_{\text{tiles}}$ into vision token embeddings $T_{vis}=[T^{lr}_{vis},T^{hr}_{vis}]$, which contain low-resolution tokens $T^{lr}_{vis}$ from the $I_{\text{thumb}}$ and higher-resolution $T^{hr}_{vis}$ from the $I_{\text{tiles}}$. $T_{vis}$ are concatenated with text token embeddings $T_{txt}$ from the tokenized text instruction, forming a multimodal token sequence $T=[T_{vis},T_{txt}]$. The LVLM processes $T$ via transformer-based decoder layers with causal self-attention, where the attention score $A$ can be represented as:
@@
Figure 3: The proposed RFM and attention distillation strategy. The left part indicates our core idea: distill accurate text-related key region localization ability from the LLM part of the LVLM. The right part shows the distillation details. We only select specific layer pairs for distillation to avoid hidden state discontinuities. “sys token" represents the tokens from the system prompt.
@@ 4.1 Method
To achieve efficient understanding of ultra-high-resolution images, our core idea is coarse-to-fine focusing: we first gain a low-resolution overview, then iteratively zoom into text-related regions for details, as shown in Fig. 2. Implementing this necessitates a multi-resolution representation. Therefore, we first introduce the Dynamic Image Pyramid (DIP) in Sec. 4.1.1. Then we need to identify where to focus next. To solve this, we propose the Region Focus Module (RFM) to provide attention distribution for input vision tokens (Sec. 4.1.2). Finally, these components are integrated into an iterative refinement process (Sec. 4.1.3).
@@ 4.1.1 The Construction of Dynamic Image Pyramid
As established, our coarse-to-fine strategy requires a multi-resolution representation of the image $I_{img}$. We realize this through the DIP, a structure designed to provide image tiles at various ground sample distances (GSDs).
First, we iteratively downsample $I_{img}$ by a factor of 2, generating a series of images $\{I^{1}_{init},I^{2}_{init},\dots,I^{P}_{init}\}$ until the shorter side reaches a pre-defined minimum length. Let a basic image tile size be $B\times B$ (e.g., $336\times 336$ for CLIP-L14), for the $p$-th scaled image $I^{p}_{init}$, we calculate the number of image tiles along the height and width as follows:
@@ 4.1.2 Attention Distillation for Key Region Focus
With the DIP established, the central challenge becomes generating a reliable guidance signal to determine where to focus next. While a sliding-window similarity map from VLMs like CLIP is a straightforward baseline, we posit that the attention maps from the LVLM’s LLM part provide a superior localization signal. This is inspired by prior work [23, 72, 81] where LLM attention tends to converge on key visual tokens. To support this, we quantitatively validate that LLM attention significantly outperforms the CLIP/RemoteCLIP baselines by offering more accurate, text-related localization signals in Tab. 7.
To harness this high-quality guidance efficiently, we introduce the Region Focus Module (RFM), a lightweight module trained via attention distillation to mimic the text-guided localization capability of the LLM teacher, as shown in Fig. 3. After distillation, the trained RFM can guide the selection of key image tiles and prune irrelevant vision tokens before the main LLM computation.
Specifically, assuming the LLM has $M$ layers $l_{1},l_{2},\dots,l_{M}$, the proposed RFM adopts the same architecture with $R$ layers $r_{1},r_{2},\dots,r_{R}$, where $R$$<$$M$. Each RFM layer $r_{i}$ is initialized from a selected LLM layer $l_{m_{i}}$, where $\{m_{1},m_{2},\dots,m_{R}\}$ forms a sparse subset of LLM layer indices. For layer-wise distillation, we select $K$ pairs ($K$$<$$R$) from RFM and LLM as student-teacher pairs, represented as $\{(r_{k},l_{m_{k}})\mid k=1,2,\dots,K\}$. This strategy avoids rigidly aligning each $r_{i}$ with $l_{m_{i}}$, which would neglect the non-teacher layers in the LLM, resulting in inconsistency between adjacent RFM layers.
@@ 4.1.3 Text-Guided Token Pruning with Pyramid
Having introduced the DIP and the RFM, this section details how they are integrated into a cohesive, iterative process. This process achieves our coarse-to-fine strategy by selectively retrieving high-resolution tiles and pruning irrelevant tokens, as illustrated in Fig. 2.
The specific inference procedure, detailed in Alg. 1, is as follows. For simplicity, the tokens from the system prompt are omitted. We initialize the vision input using $I_{\text{thumb}}$ and $I^{p}_{\text{tiles}}$ to generate the initial $T_{vis}=[T^{lr}_{vis},T^{p,hr}_{vis}]$, where $p=1$, then the concatenated tokens $T$ are fed into the RFM. From the last layer of the RFM, we compute the average attention scores from all heads in the MHSA, and extract the attention scores $\smash{A^{K,hr}_{\text{stu}}}$ of $\smash{T^{p,hr}_{\text{vis}}}$. By selecting the top-$\lambda$ proportion from $\smash{A^{K,hr}_{\text{stu}}}$, we identify the indices of key tokens and map their coordinates to image tile-level coordinates, to select key image tiles $I^{p+1}_{\text{key}}$ from $I^{p+1}_{\text{tiles}}$. Based on the number of $I^{p+1}_{\text{key}}$, we determine whether to prune the $\smash{T^{p,hr}_{\text{vis}}}$ directly or replace it with vision tokens $\smash{T^{p+1,hr}_{\text{vis}}}$ from higher-resolution tiles. If $\smash{T^{p+1,hr}_{\text{vis}}}$ is obtained, the tile selection process is repeated recursively until the last layer of DIP is reached. This approach enables LVLM to focus only on processing a few high-resolution image tiles in a coarse-to-fine manner, thereby reducing computational complexity while preserving critical text-related image details.
@@ 4.2 The Construction of LRS-VQA Benchmark
Considering the limitations of existing benchmarks for evaluating LVLMs’ perception of RSIs in terms of question diversity and image size, we propose a new benchmark for more comprehensive evaluation.
@@ 4.2.1 Annotation Pipeline
The annotation pipeline of LRS-VQA is illustrated in Fig. 4. We collect 3 publicly remote sensing datasets: FAIR1M-1.0 [62], GLH-Bridge [35], and STAR [37], and identify all unique targets based on the object detection labels to generate unique references from them (e.g., “the top-most ship" and “the largest dock"). Subsequently, we crop the region with the unique target from the original image, and feed it, along with a prompt containing the unique reference, into GPT-4V [54] to generate question-answer pairs about the target’s color, shape, status, etc. Details of the overall construction process are provided in the Appendix A.
@@ 5 Experiments
w/ PruMerge++ [58] 43.27 28.38 32.78 34.86 0.152
w/ VisionZip [77] 42.71 24.55 34.55 33.98 0.183
w/ FastV [8] 39.52 30.10 43.99 37.93 0.192
w/ PDrop [72] 41.00 30.54 47.77 39.85 0.184
w/ prune (CLIP) 37.80 27.65 44.90 36.86 0.171
w/ prune (RemoteCLIP) 39.40 29.83 47.45 38.77 0.148
@@
Table 3: Comparison with different token reduction methods based on LLaVA-Next-Qwen2. The “anyres-p25” setting means the maximum number of pre-defined grids is 25 (1680×1680 pixels). FPS is computed based on the inference time for continuous 800 images from MME-Realworld-RS.
@@
w/ PruMerge++ [58] 83520 5328 43.75
w/ VisionZip [77] 83520 9280 83.56
w/ FastV [8] 83520 21312 109.21
w/ PDrop [72] 83520 21312 101.62
@@ 5.1 Experimental Details
Data. Our pre-training (PT) phase is the same as LLaVA-1.5 [42] (using 558K data). For the SFT phase, we collect 484K samples, including 300K sampled from LLaVA-1.5-665K, 146K sampled from RSVQA-HR [47], and 38K template-based samples from three RS datasets [15, 35, 37] using their labels. Within this 38k data, some images share the same source with LRS-VQA, but there is no overlap between them. Details can be found in the Appendix B.1.
Benchmarks and evaluation metrics: i) MME-RealWorld-RS: the RS part of MME-Realworld [91], containing 1,298 RSIs with expert-annotated questions in three types: color, count, and position. We follow the official evaluation script but modify the prompt by removing “The best answer is:" to address Vicuna-1.5-based models’ tendency to respond with only option A. ii) LRS-VQA: it consists of 3 parts: LRS-FAIR, LRS-Bridge, and LRS-STAR, containing 2,272, 1,062, and 3,999 QA pairs, respectively. For the short open-ended format, we adopt a structured evaluation metric following [22, 71], using WordNet [52], with a semantic similarity threshold of 0.8.
Experimental settings. We evaluate existing LVLMs using their official weights or APIs. For fair comparison, we also re-implement some methods on our 484k SFT dataset. Our approach is built upon the LLaVA-1.5 codebase, where we integrate the anyres strategy and Qwen2 language model. All experiments are conducted on 4 A100 80GB GPUs. For our method, the minimum length for DIP is 1,008 pixels. The $N_{\max}$ is set to 40 and 80 for Vicuna-1.5 and Qwen2, respectively. The vision token saving ratio $\alpha$ is 0.25, $\lambda_{\text{hr}},\lambda_{\text{mse}},\lambda_{\text{kl}}$ are set to 2.0, 1.0, 1.0. The RFM-LLM layer pairs are $[1,5,11,14]$ with distillation applied to the first and last pairs. Additional experimental details are provided in Appendix B.2.